What is the mass of 4.00x10^20 CO2 molecules
1
Expert's answer
2019-10-15T07:45:01-0400

Solution.

"\frac{m}{M} = \frac{N}{Na}"

"m = \frac{M \times N}{Na}"

"m = \frac{44 \times 4*10^{20}}{6.02*10^{23}} = 0.03 \ g"

m = 0.03 g

Answer:

m = 0.03 g